# The adjacent Hindman’s theorem and the $\mathbb{Z}$-Ramsey theorem

### v1

@@ Abstract.
We consider the weakening of Ramsey’s theorem that arises from considering only translation-invariant colourings of pairs, and show that this has the same strength (both from the viewpoint of Reverse Mathematics and from the viewpoint of Computability Theory) as the adjacent Hindman’s theorem, proposed by L. Carlucci (Arch. Math. Log. 57 (2018), 381–359). We also investigate some higher dimensional versions of both of these statements.
@@ 1. Introduction
In this paper, we consider some Ramsey-theoretic combinatorial results from the perspective of both Computability Theory and Reverse Mathematics. Ramsey’s theorem on dimension $n$, denoted $\mathnormal{\mathsf{RT}}^{n}$ ($n\geq 1$), is the statement that for every $c:[\mathbb{N}]^{n}\longrightarrow k$ (where $k$ is an arbitrary finite number), there exists an infinite set $X\subseteq\mathbb{N}$ such that $[X]^{n}$ is $c$-monochromatic. On the other hand, Hindman’s theorem, denoted $\mathnormal{\mathsf{HT}}$, is the statement that for every $c:\mathbb{N}\longrightarrow k$ there exists an infinite set $X\subseteq\mathbb{N}$ such that the set
@@
These two principles have been extensively studied, and constitute an important vein of contemporary research in the interface between combinatorics and (various branches of) logic. For example, it is known that, if $n\geq 3$, then $\mathnormal{\mathsf{RT}}^{n}$ is equivalent to $\mathnormal{\mathsf{ACA}_{0}}$ over $\mathnormal{\mathsf{RCA}_{0}}$; somewhat surprisingly, $\mathnormal{\mathsf{RT}}^{2}$ is a principle at the same time strictly weaker than $\mathnormal{\mathsf{ACA}_{0}}$ and strictly stronger than $\mathnormal{\mathsf{WKL}_{0}}$, whereas $\mathnormal{\mathsf{RT}}^{1}$ is simply the pigeonhole principle (strictly stronger than plain $\mathnormal{\mathsf{RCA}_{0}}$ but much weaker than $\mathnormal{\mathsf{WKL}_{0}}$). On the other hand, it is also known that, again over $\mathnormal{\mathsf{RCA}_{0}}$, $\mathnormal{\mathsf{ACA}_{0}^{+}}$ implies $\mathnormal{\mathsf{HT}}$ which in turn implies $\mathnormal{\mathsf{ACA}_{0}}$, where $\mathnormal{\mathsf{ACA}_{0}^{+}}$ is essentially $\mathnormal{\mathsf{ACA}_{0}}$ plus the existence of the $\omega$-th Turing jump of every set. The precise strength of $\mathnormal{\mathsf{HT}}$ in the hierarchy of reverse-mathematical principles is still unknown, and determining this strength is currently one of the important open problems in Reverse Mathematics.
@@
In this paper, we consider certain weakenings of each of the two principles mentioned above. On the front of Hindman’s theorem, we shall work with L. Carlucci’s Adjacent Hindman’s Theorem proposed in [2], which results from weakening the monochromaticity requirement not to a full set $\operatorname{FS}(X)$ but rather to the restricted set $\operatorname{AFS}(X)$ of adjacent finite sums, meaning those sums of finitely many consecutive elements of $X$ when ordered increasingly, leaving no gaps. Carlucci showed that, over $\mathnormal{\mathsf{RCA}_{0}}$, $\mathnormal{\mathsf{RT}}^{2}$ implies $\mathnormal{\mathsf{AHT}}$, even with an extra condition that we will also consider, called apartness. On the other hand, we will consider a weakening of Ramsey’s theorem that arises from restricting the allowable problems for which we require a solution —concretely, stating the theorem only for those colourings that are invariant under the translation action of $\mathbb{Z}$, obtaining what we will call the $\mathbb{Z}$-Ramsey’s theorem. Although this principle has not been previously studied within Reverse Mathematics, the idea of restricting instances of Ramsey’s theorem only to translation-invariant colourings has already been used, e.g., by Petrenko and Protasov in their definition of $\mathbb{Z}$-Ramsey ultrafilters [9].
@@
Many of the principles mentioned in the previous paragraph turn out to be equivalent to $\mathnormal{\mathsf{ACA}_{0}}$ over $\mathnormal{\mathsf{RCA}_{0}}$, so for the most part the viewpoint of Reverse Mathematics will not be very informative in that respect. In view of this, we focus more on the Computability Theoretic aspect of the principles, thinking instead about the reducibility relation by means of which an instance of one problem can be turned into an instance of a different problem and a solution to the latter gets turned back into a solution to the original instance. More concretely, we will work with uniform reducibility, sometimes also known as (or a particular case of) Weihrauch reducibility [^1] In fact, the proofs of the theorems in this paper actually establish what is known as a strong Weihrauch reducibility, since e.g. the Turing functionals for turning a solution to one problem into a solution of the other do not require information from the concrete instance of the problem at hand (more concretely, in the notation explained below we have that $\Psi^{X}$, rather than $\Psi^{X\oplus Y}$, is already a solution to $Y$). We are thankful to L. Carlucci for this observation.. Recall [7] that a problem $P$ is uniformly reducible to another problem $Q$, denoted $P\leq_{u}Q$, if there are Turing functionals $\Phi,\Psi$ such that for every instance $X$ of $P$, $\Phi^{X}$ is an instance of $Q$, and for every solution $Y$ to $\Phi^{X}$, $\Psi^{X\oplus Y}$ is a solution to $X$. So we will use this notion of reducibility to gauge the position of these principles within the computability-theoretic hierarchy; all of our proofs can be turned, by simply looking at them carefully enough, into proofs in $\mathnormal{\mathsf{RCA}_{0}}$, thus yielding corresponding (but coarser) Reverse Mathematics results as corollaries.
@@ Acknowledgements
We thank Lorenzo Carlucci for useful discussions and suggestions. The second author was partially supported by internal grants SIP-20221862 and SIP-20240886 from the Instituto Politécnico Nacional.
2. The $\mathbb{Z}$-invariant Ramsey theorem
@@
    Given an $n\in\mathbb{N}$, we will say that a colouring of $n$-tuples $c:[\mathbb{N}]^{n}\longrightarrow k$ is $\mathbb{Z}$-invariant if, for every $n$-tuple $\{x_{1},\ldots,x_{n}\}\in[\mathbb{N}]^{n}$ and for every $z\in\mathbb{N}$, it is the case that $c(\{x_{1},\ldots,x_{n}\})=c(\{x_{1}+z,\ldots,x_{n}+z\})$.
@@
    The $\mathbb{Z}$-Ramsey theorem for $n$-tuples, which we will denote by -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}$, is the statement that every $\mathbb{Z}$-invariant colouring of $n$-tuples admits an infinite monochromatic set.
It is clear that, for each $n\in\mathbb{N}$, -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}$ is a weaker statement than $\mathnormal{\mathsf{RT}}^{n}$. The following theorem basically establishes that -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n+1}$ implies $\mathnormal{\mathsf{RT}}^{n}$ over $\mathnormal{\mathsf{RCA}_{0}}$, although, as we warned in the introduction, we will attempt to provide more precise information by phrasing our theorem in the language of uniform reducibility.
@@
For each $n\geq 2$, $\mathnormal{\mathsf{RT}}^{n}\leq_{u}\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n+1}$.
@@
where $x_{0}=\min(A)$. If $[A]^{n+1}$ is $\Phi^{c}$-monochromatic (say, in colour $i$) then $[\Psi^{A}]^{n}$ is $c$-monochromatic (in the same colour $i$), since for any elements $x_{1}-x_{0},\ldots,x_{n}-x_{0}\in\Psi^{A}$ we have
@@
Since the statements $\mathnormal{\mathsf{RT}}^{n}$ are equivalent for $n\geq 3$ (and they are equivalent to $\mathnormal{\mathsf{ACA}_{0}}$), the only interesting corollary, from the perspective of Reverse Mathematics, is the part where $n\leq 3$.
@@
    -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}$ is equivalent to $\mathnormal{\mathsf{ACA}_{0}}$ for each $n\geq 4$, and
@@
    $\mathnormal{\mathsf{ACA}_{0}}\Rightarrow\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{3}\Rightarrow\mathnormal{\mathsf{RT}}^{2}\Rightarrow\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}\Rightarrow\mathnormal{\mathsf{RT}}^{1}$.
Later on, in the last section (and once we have more tools under our belt), we come back to whether the first two arrows from part (2) of Corollary 3 are reversible.
@@ 2.1. Relation with the Adjacent Hindman’s Theorem
is monochromatic, where the sequence $(x_{n}\big{|}n\in\mathbb{N})$ represents the (unique) increasing enumeration of the set $X$.
Given that the set $\operatorname{AFS}(X)$ is defined in such a way that the order within our set matters (unlike the usual $\operatorname{FS}$ sets utilized in Hindman’s theorem), one could conceivably state a version of the $\mathnormal{\mathsf{AHT}}$ for sequences rather than sets, i.e. affirming the existence of a sequence $(x_{n}\big{|}n\in\mathbb{N})$, not necessarily increasing (or even injective!) whose adjacent finite sums are monochromatic. However, it is readily seen that these two versions of the $\mathnormal{\mathsf{AHT}}$ must be equivalent (whether from the viewpoint of Reverse Mathematics over $\mathnormal{\mathsf{RCA}_{0}}$, or from the viewpoint of Computability Theory). Clearly the original version as stated by Carlucci implies the “sequence” version (by replacing the set $X$ with the sequence that increasingly enumerates its elements); conversely, given an arbitrary sequence $x_{n}$, one can recursively define $y_{1}=x_{1}$, $k_{1}=1$; and $y_{n+1}=x_{k_{n}+1}+x_{k_{n}+2}+\cdots+x_{k_{n+1}}$, with $k_{n+1}$ the least number making the $y_{n+1}$ thus defined to be strictly larger than $y_{n}$. This way we obtain a strictly increasing sequence $(y_{n}\big{|}n\in\mathbb{N})$ (in particular, we obtain a set $Y$ whose increasing enumeration is precisely the sequence of $y_{n}$) such that $\operatorname{AFS}(Y)=\operatorname{AFS}(y_{n}\big{|}n\in\mathbb{N})\subseteq\operatorname{AFS}(x_{n}\big{|}n\in\mathbb{N})$.
@@
-$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}\leq_{u}\mathnormal{\mathsf{AHT}}$ and $\mathnormal{\mathsf{AHT}}\leq_{u}\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}$.
@@
Since numbers of the form $y_{n+1}+\cdots+y_{m}$ are precisely the elements of $\operatorname{AFS}(Y)$, the conclusion is that $\operatorname{AFS}(Y)$ is $\Phi_{1}^{c}$-monochromatic if and only if $[\Psi_{1}^{Y}]^{2}$ is $c$-monochromatic.
@@
Note that the previous proof, in a sense, shows that $\mathbb{Z}$-invariant colourings are precisely those colourings that depend only on the distance between the two elements of the pair being coloured. Of course, we now have the following corollary, whose proof consists of carefully carrying out the previous proof within Recursive Arithmetic.
@@
Over $\mathnormal{\mathsf{RCA}_{0}}$, $\operatorname{AFS}$ is equivalent to -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}$.
@@
We now proceed to study a higher-dimensional version of the Adjacent Hindman’s Theorem, much in the same spirit that the Milliken–Taylor theorem constitutes a two-dimensional version of the usual Hindman’s theorem. Recall that the Milliken–Taylor theorem ensures that, given any colouring of pairs of finite subsets of $\mathbb{N}$, there exists a pairwise disjoint family such that all ordered pairs of finite unions have the same colour. In the adjacent context, we will have to consider adjacent pairs of adjacent finite unions, and similarly for higher dimensions.
@@
For each $d\in\mathbb{N}\setminus\{0\}$, we have -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{d+1}\leq_{u}\mathnormal{\mathsf{AHT}}^{d}$ and $\mathnormal{\mathsf{AHT}}^{d}\leq_{u}\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{d+1}$.
@@
The Turing functionals $\Phi_{1},\Psi_{1}$ witnessing that -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{d+1}\leq_{u}\mathnormal{\mathsf{AHT}}^{d}$ are: for a $\mathbb{Z}$-invariant colouring $c:[\mathbb{N}]^{d+1}\longrightarrow k$, we let $\Phi_{1}^{c}:\mathbb{N}^{d}\longrightarrow k$ be given by
@@
where we have defined $z_{i}=x_{i}-x_{i-1}=y_{k_{i-1}+1}+\cdots+y_{k_{i}}$. So the values of $c$ on $(d+1)$-tuples from $\Psi_{1}^{Y}$ are exactly the values of $\Phi_{1}^{c}$ on $d$-tuples from $Y$ that have the form

### v2

# The adjacent Hindman’s theorem and the $\mathbb{Z}$-Ramsey’s theorem

@@ Abstract.
We consider the restriction of Ramsey’s theorem that arises from considering only translation-invariant colourings of pairs, and show that this has the same strength (both from the viewpoint of Reverse Mathematics and from the viewpoint of Computability Theory) as the Adjacent Hindman’s Theorem, proposed by L. Carlucci (Arch. Math. Log. 57 (2018), 381–359). We also investigate some higher dimensional versions of both of these statements.
Key words and phrases:
Ramsey-type theorem, Ramsey’s theorem, Hindman’s theorem, Computability theory, Reverse mathematics.
2020 Mathematics Subject Classification:
Primary 03F35, 03D30, Secondary 05D10.
@@ 1. Introduction
In this paper, we consider some Ramsey-theoretic combinatorial results from the perspective of both Computability Theory and Reverse Mathematics. Ramsey’s theorem in dimension $n$, denoted $\mathnormal{\mathsf{RT}}^{n}$ ($n\geq 1$), is the statement that for every $c:[\mathbb{N}]^{n}\longrightarrow k$ (where $k$, an arbitrary finite number, is used to denote the set $\{0,\ldots,k-1\}$), there exists an infinite set $X\subseteq\mathbb{N}$ such that $X]^{n}$ is $c$-monochromatic. On the other hand, Hindman’s theorem, denoted $\mathnormal{\mathsf{HT}}$, is the statement that for every $c:\mathbb{N}\longrightarrow k$ there exists an infinite set $X\subseteq\mathbb{N}$ such that the set
@@
These two principles have been extensively studied, and constitute an important vein of contemporary research in the interface between combinatorics and (various branches of) logic. In order to mention some previous results, we proceed to briefly describe the systems involved. $\mathnormal{\mathsf{RCA}}_{0}$ is a subsystem of Second-Order Arithmetic that, roughly speaking, has Peano Arithmetic as its first-order part, and includes closure under relative computability as its second-order part. $\mathnormal{\mathsf{WKL}}_{0}$ is $\mathnormal{\mathsf{RCA}}_{0}$ plus closure under taking an infinite branch of each binary tree (i.e. the statement of König’s Lemma, restricted to binary trees, holds), and $\mathnormal{\mathsf{ACA}}_{0}$ is $\mathnormal{\mathsf{RCA}}_{0}$ plus closure under taking Turing jumps. It is well-known that $\mathnormal{\mathsf{ACA}}_{0}$ is strictly stronger than $\mathnormal{\mathsf{WKL}}_{0}$, which in turn is strictly stronger than $\mathnormal{\mathsf{RCA}}_{0}$. The reader might consult [16] for a thorough introduction to this area of research, or [11] for a treatment more readily oriented to the three systems that we just mentioned and their relationship to Ramsey theory in general. Nowadays, it is known that, if $n\geq 3$, then $\mathnormal{\mathsf{RT}}^{n}$ is equivalent to $\mathnormal{\mathsf{ACA}}_{0}$ over $\mathnormal{\mathsf{RCA}}_{0}$; somewhat surprisingly, $\mathnormal{\mathsf{RT}}^{2}$ is a principle at the same time strictly weaker than $\mathnormal{\mathsf{ACA}}_{0}$ and incomparable with than $\mathnormal{\mathsf{WKL}}_{0}$, whereas $\mathnormal{\mathsf{RT}}^{1}$ is simply the pigeonhole principle (strictly stronger than plain $\mathnormal{\mathsf{RCA}}_{0}$ but much weaker than $\mathnormal{\mathsf{WKL}}_{0}$). These results are classical, although the interested reader can consult, e.g., [8] (especially Corollaries 8.2.6 and 8.6.2) for a more contemporary source. On the other hand, it is also known that, again over $\mathnormal{\mathsf{RCA}}_{0}$, $\mathnormal{\mathsf{ACA}}_{0}^{+}$ implies $\mathnormal{\mathsf{HT}}$ which in turn implies $\mathnormal{\mathsf{ACA}}_{0}$, where $\mathnormal{\mathsf{ACA}}_{0}^{+}$ is essentially $\mathnormal{\mathsf{ACA}}_{0}$ plus closure under the $\omega$-th Turing jump; these last results are due to Blass, Hirst and Simpson [1]. The precise strength of $\mathnormal{\mathsf{HT}}$ in the hierarchy of reverse-mathematical principles is still unknown, and determining this strength is currently one of the important open problems in Reverse Mathematics. A good overview of the state of the art regarding this problem can be found in [3].
@@
In this paper, we consider certain restrictions of Hindman’s theorem and of Ramsey’s theorem. On the front of Hindman’s theorem, we shall work with L. Carlucci’s Adjacent Hindman’s Theorem proposed in [2], which results from restricting the monochromaticity requirement not to a full set $\operatorname{FS}(X)$ but rather to the restricted set $\operatorname{AFS}(X)$ of adjacent finite sums, meaning those sums of finitely many consecutive elements of $X$ when ordered increasingly, leaving no gaps. Carlucci showed that, over $\mathnormal{\mathsf{RCA}}_{0}$, $\mathnormal{\mathsf{RT}}^{2}$ implies $\mathnormal{\mathsf{AHT}}$, even with an extra condition that we will also consider, called apartness. On the other hand, we will consider a restriction of Ramsey’s theorem that arises from restricting the allowable problems for which we require a solution —concretely, stating the theorem only for those colourings that are invariant under the translation action of $\mathbb{Z}$, obtaining what we will call the $\mathbb{Z}$-Ramsey’s theorem. Although this principle has not been previously studied within Reverse Mathematics, the idea of restricting instances of Ramsey’s theorem only to translation-invariant colourings has already been used, e.g., by Petrenko and Protasov in their definition of $\mathbb{Z}$-Ramsey ultrafilters [15].
@@
Many of the principles mentioned in the previous paragraph turn out to be equivalent to $\mathnormal{\mathsf{ACA}}_{0}$ over $\mathnormal{\mathsf{RCA}}_{0}$, so for the most part the viewpoint of Reverse Mathematics will not be very informative in that respect. In view of this, we focus more on the Computability-theoretic aspect of the principles, thinking instead about the reducibility relation, by means of which an instance of one problem can be turned into an instance of a different problem and a solution to the latter gets turned back into a solution to the original instance. More concretely, we will work with Weihrauch reducibility [^1] Some authors (e.g., [11, p. 22]) have called this notion uniform reducibility, but the name “Weihrauch reducibility” seems to be the most commonly used nowadays. We are grateful to two anonymous referees for this observation.. Recall [8, Definition 4.3.1] that a problem $P$ is Weihrauch reducible to another problem $Q$, denoted $P\mathop{\leq_{\mathsf{W}}}Q$, if there are Turing functionals $\Phi,\Psi$ such that for every instance $X$ of $P$, $\Phi^{X}$ is an instance of $Q$, and for every solution $Y$ to $\Phi^{X}$, $\Psi^{X\oplus Y}$ is a solution to $X$. If we have the exact same situation, but with a Turing functional $\Psi$ such that $\Psi^{Y}$ is a solution to $X$ (i.e. we may compute the solution to $X$ just with information from $Y$ without needing to use any information about $X$), then we say that $P$ is strongly Weihrauch reducible [8, Definition 4.4.1] to $Q$ and write $P\mathop{\leq_{\mathrm{s}\mathsf{W}}}Q$. If each of the problems $P,Q$ is Weihrauch reducible (respectively, strongly Weihrauch reducible) to the other, then we say that $P$ is Weihrauch equivalent (respectively, strongly Weihrauch equivalent) to $Q$ and we denote this by $P\mathop{\equiv_{\mathsf{W}}}Q$ (respectively, $P\mathop{\equiv_{\mathrm{s}\mathsf{W}}}Q$). So we will use this notion of reducibility to gauge the position of the Adjacent Hindman’s Theorem and the $\mathbb{Z}$-Ramsey’s Theorem within the computability-theoretic hierarchy; an important feature of all of our proofs is that they can be formalized within $\mathnormal{\mathsf{RCA}}_{0}$, thus yielding corresponding (but coarser) Reverse Mathematics results as corollaries.
@@ Acknowledgements
We thank Lorenzo Carlucci for useful discussions and suggestions (including the realization that our proofs yielded not only Weihrauch reductions, but strong Weihrauch reductions), as well as two anonymous referees whose thorough and abundant comments helped greatly improve this paper. The second author was partially supported by IPN’s internal grants SIP-20221862, SIP-20240886, and SIP-20253559, as well as by Secihti’s grant CBF2023-2024-334. The first and fourth authors were supported by scholarships from Secihti.
2. The $\mathbb{Z}$-invariant Ramsey’s theorem
@@
    Given an $n\geq 1$, we will say that a colouring of $n$-subsets $c:[\mathbb{N}]^{n}\longrightarrow k$ is $\mathbb{Z}$-invariant if, for every $n$-subset $\{x_{1},\ldots,x_{n}\}\in[\mathbb{N}]^{n}$ and for every $z\in\mathbb{N}$, it is the case that $c(\{x_{1},\ldots,x_{n}\})=c(\{x_{1}+z,\ldots,x_{n}+z\})$.
@@
    The $\mathbb{Z}$-Ramsey’s theorem for $n$-subsets, which we will denote by -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}$, is the statement that every $\mathbb{Z}$-invariant colouring of $n$-subsets admits an infinite monochromatic set.
It is clear that, for each $n\geq 1$, -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}$ is a trivial corollary of $\mathnormal{\mathsf{RT}}^{n}$. Now, in the case of $\mathnormal{\mathsf{RT}}^{n}$, there is no difference (whether from the viewpoint of Weihrauch equivalence, or over $\mathnormal{\mathsf{RCA}}_{0}$) whether one formulates it over $\mathbb{N}$ or over any infinite $X\subseteq\mathbb{N}$ [8, Proposition 8.3.4]. This is not the case for -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}$: for example, if $X\subseteq\mathbb{N}$ is any infinite set with the property that (if we enumerate its elements increasingly by $x_{n}$) $x_{n+1}-x_{n}>x_{n}-x_{n-1}$ for all $n$, then (since no two distinct pairs of elements of $X$ are at the same distance from one another) the family of $\mathbb{Z}$-invariant colourings restricted to $X$ is the same as the family of all arbitrary colourings on $X$. Therefore, for such a set, the version of -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}$ stated for $X$ would be Weihrauch-equivalent to the full $\mathnormal{\mathsf{RT}}^{n}$. We therefore refrain from considering “general” versions of the $\mathbb{Z}$-Ramsey theorem.
The following theorem basically establishes that -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n+1}$ implies $\mathnormal{\mathsf{RT}}^{n}$ over $\mathnormal{\mathsf{RCA}}_{0}$, although, as we warned in the introduction, we provide more precise information by phrasing our theorem in the language of Weihrauch reducibility.
@@
For each $n\geq 1$, $\mathnormal{\mathsf{RT}}^{n}\mathop{\leq_{\mathrm{s}\mathsf{W}}}\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n+1}$.
@@
where $x_{0}=\min(A)$. If $A]^{n+1}$ is $\Phi^{c}$-monochromatic (say, in colour $i$) then $\Psi^{A}]^{n}$ is $c$-monochromatic (in the same colour $i$), since for any elements $x_{1}-x_{0},\ldots,x_{n}-x_{0}\in\Psi^{A}$ we have
@@
The attentive reader should note that, in the previous proof, in order for $\Psi^{A}]^{n}$ to be $c$-monochromatic, we do not fully use the fact that all $(n+1)$-subsets of $A$ are $\Psi^{c}$-monochromatic, but only those whose first element is $x_{0}$—for example, for $n=1$, we only need the $2$-subsets taking their first element from $\{x_{0}\}$ and their second element from $\{x_{n}\big|n\geq 1\}$, which looks like a very particular case of a Polarized version of Ramsey’s Theorem. This suggests an intriguing question for possible further research, namely whether some restriction of the Increasing Polarized Ramsey’s Theorem, possibly in a higher dimension, can already imply Ramsey’s Theorem—and what would be the right restriction to make in order to obtain an equivalence.
The proof of Theorem 2 is simple enough that, by carefully going through it from the perspective of Reverse Mathematics, it is apparent that no extra induction beyond $\mathnormal{\mathsf{I}\mathrm{\Sigma}_{1}^{0}}$ is used. Hence, this proof also yields implications over $\mathnormal{\mathsf{RCA}}_{0}$, which we will summarize in the following corollary. Since the statements $\mathnormal{\mathsf{RT}}^{n}$ are equivalent for $n\geq 3$ (and they are equivalent to $\mathnormal{\mathsf{ACA}}_{0}$), the only interesting corollary, from the perspective of Reverse Mathematics, is the part where $n\leq 3$.
@@
    For each $n\geq 4$, $\mathnormal{\mathsf{RCA}}_{0}\vdash\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}\leftrightarrow\mathnormal{\mathsf{ACA}}_{0}$,
@@
    $\mathnormal{\mathsf{ACA}}_{0}\vdash\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{3}$,
  3. (3)
    $\mathnormal{\mathsf{RCA}}_{0}\vdash\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{3}\rightarrow\mathnormal{\mathsf{RT}}^{2}\rightarrow\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}\rightarrow\mathnormal{\mathsf{RT}}^{1}$.
A priori, the arrows on part (3) of Corollary 3 may or may not be reversible. Later on, in the last section (and once we have more tools under our belt), we will look into this specific question regarding the arrow from -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{3}$ to $\mathnormal{\mathsf{RT}}^{2}$.
@@ 2.1. Relation with the Adjacent Hindman’s Theorem
is monochromatic, where the sequence $\langle x_{n}\big|n\in\mathbb{N}\rangle$ represents the (unique) increasing enumeration of the set $X$.
Given that the set $\operatorname{AFS}(X)$ is defined in such a way that the order within our set matters (unlike the usual $\operatorname{FS}$ sets utilized in Hindman’s theorem), one could conceivably state a version of the $\mathnormal{\mathsf{AHT}}$ for sequences rather than sets, i.e. affirming the existence of a sequence $\langle x_{n}\big|n\in\mathbb{N}\rangle$, not necessarily increasing (or even injective!) whose adjacent finite sums are monochromatic. However, it is readily seen that these two versions of the $\mathnormal{\mathsf{AHT}}$ must be equivalent (whether from the viewpoint of Reverse Mathematics over $\mathnormal{\mathsf{RCA}}_{0}$, or from the viewpoint of Computability Theory). Clearly the original version as stated by Carlucci implies the “sequence” version (by replacing the set $X$ with the sequence that increasingly enumerates its elements); conversely, given an arbitrary sequence $x_{n}$, one can recursively define $y_{1}=x_{1}$, $k_{1}=1$; and $y_{n+1}=x_{k_{n}+1}+x_{k_{n}+2}+\cdots+x_{k_{n+1}}$, with $k_{n+1}$ the least number making the $y_{n+1}$ thus defined to be strictly larger than $y_{n}$. This way we obtain a strictly increasing sequence $\langle y_{n}\big|n\in\mathbb{N}\rangle$ (in particular, we obtain a set $Y$ whose increasing enumeration is precisely the sequence of $y_{n}$) such that $\operatorname{AFS}(Y)=\operatorname{AFS}(y_{n}\big|n\in\mathbb{N})\subseteq\operatorname{AFS}(x_{n}\big|n\in\mathbb{N})$.
@@
-$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}\mathop{\equiv_{\mathrm{s}\mathsf{W}}}\mathnormal{\mathsf{AHT}}$.
@@
Since numbers of the form $y_{n+1}+\cdots+y_{m}$ are precisely the elements of $\operatorname{AFS}(Y)$, the conclusion is that $\operatorname{AFS}(Y)$ is $\Phi_{1}^{c}$-monochromatic if and only if $\Psi_{1}^{Y}]^{2}$ is $c$-monochromatic.
@@
Note that the previous proof, in a sense, shows that $\mathbb{Z}$-invariant colourings are precisely those colourings that depend only on the distance between the two elements of the pair being coloured. Once again, the proof is simple enough that one can directly see that no induction beyond $\mathnormal{\mathsf{I}\mathrm{\Sigma}_{1}^{0}}$ is used; therefore, the same proof (appropriately formalized) yields as a result the following corollary.
@@
$\mathnormal{\mathsf{RCA}}_{0}\vdash\mathnormal{\mathsf{AHT}}\leftrightarrow\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}$.
@@
We now proceed to study a higher-dimensional version of the Adjacent Hindman’s Theorem, much in the same spirit that the Milliken–Taylor theorem constitutes a higher-dimensional version of the usual Hindman’s theorem. Recall first that Hindman’s theorem has an equivalent formulation in terms of finite unions: for every colouring of the set of finite subsets of $\mathbb{N}$, there exists an infinite, pairwise disjoint family of such finite subsets satisfying that all of the unions of finitely many of those sets receive the same colour. The Milliken–Taylor theorem [14, 18], for example for dimension $2$, ensures that, given any colouring of pairs of finite subsets of $\mathbb{N}$, there exists an infinite pairwise disjoint family such that all ordered pairs of finite unions have the same colour [^2] A pair of finite sets $(a,b)$ is ordered if $\max(a)<\min(b)$. We switched from the finite-sums to the finite-unions formulation here in order to be able to use this notion; the Milliken–Taylor theorem can also be formulated in terms of finite-sums but then one must introduce the $\lambda$ and $\mu$ functions, which we will only do in the next subsection.; and analogously for higher-dimensional versions. In the adjacent context, we will have to consider adjacent pairs of adjacent finite unions, and similarly for higher dimensions.
@@
For each $d\in\mathbb{N}\setminus\{0\}$, we have -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{d+1}\mathop{\equiv_{\mathrm{s}\mathsf{W}}}\mathnormal{\mathsf{AHT}}^{d}$.
@@
The Turing functionals $\Phi_{1},\Psi_{1}$ witnessing that -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{d+1}\mathop{\leq_{\mathrm{s}\mathsf{W}}}\mathnormal{\mathsf{AHT}}^{d}$ are: for a $\mathbb{Z}$-invariant colouring $c:[\mathbb{N}]^{d+1}\longrightarrow k$, we let $\Phi_{1}^{c}:\mathbb{N}^{d}\longrightarrow k$ be given by
@@
where we have defined $z_{i}=x_{i}-x_{i-1}=y_{k_{i-1}+1}+\cdots+y_{k_{i}}$. So the values of $c$ on $(d+1)$-subsets from $\Psi_{1}^{Y}$ are exactly the values of $\Phi_{1}^{c}$ on $d$-tuples from $Y$ that have the form